OpenStudy (anonymous):

Using complete sentences, explain how to factor each one. 10x^2 − 7x − 3 using the ax^2+bx+c method.

OpenStudy (anonymous):

divide thro' by 10 to get x^2 -(7/10)x-3/10=0 put x terms on left numbers on right 1/2 co-eff of x term and square it(7/20)^2 and add it to both sides the left side will be a perfect square (x-7/20)^2 square root both sides and simplify (x-7/20) =3/10 +(49/400) x =1 or (-3/10)

OpenStudy (anonymous):

This is algebra. I have to make the equation. I am factoring trinomials.

OpenStudy (anonymous):

No solving for x.

OpenStudy (anonymous):

you have been asked to factor

OpenStudy (anonymous):

this can factor without all this (x-1)(10x+3) same answer

OpenStudy (anonymous):

but you are asked to use sentences

OpenStudy (anonymous):

Oh ok I got it! Thanks. I was just having a problem with the ending result!